# Learning to See and Act: Task-Aware View Planning for Robotic Manipulation

### v1

@@ Abstract
Recent vision-language-action (VLA) models for multi-task robotic manipulation commonly rely on static viewpoints and shared visual encoders, which limit 3D perception and cause task interference, hindering robustness and generalization. In this work, we propose Task-Aware View Planning (TAVP), a framework designed to overcome these challenges by integrating active view planning with task-specific representation learning. TAVP employs an efficient exploration policy, accelerated by a novel pseudo-environment, to actively acquire informative views. Furthermore, we introduce a Mixture-of-Experts (MoE) visual encoder to disentangle features across different tasks, boosting both representation fidelity and task generalization. By learning to see the world in a task-aware way, TAVP generates more complete and discriminative visual representations, demonstrating significantly enhanced action prediction across a wide array of manipulation challenges. Extensive experiments on RLBench tasks show that our proposed TAVP model achieves superior performance over state-of-the-art fixed-view approaches. Visual results and code are provided at: TAVP.
Introduction
A general-purpose robotic system is expected to accurately perform a wide range of manipulation tasks [7]. With the rapid development of vision-language-action (VLA) models [39, 119, 73, 94, 74, 111, 84, 115, 116, 117, 38, 118] and transformer-based architectures [122, 119, 73, 31, 30], many end-to-end robotic manipulation frameworks (e.g., OpenVLA[39], $\pi_{0}$[120]) have achieved promising results on complex and fine-grained tasks. Typically, these models rely on observations captured from a single view [39, 30, 42] or a few fixed viewpoints [158, 157] to guide action prediction. While this may suffice in simple environments with only a few objects, it becomes problematic in cluttered or dynamic scenes—particularly when objects or the end-effector are in motion during task execution.
@@ Introduction
Figure 1: Motivation Illustration. Observations captured from fixed cameras often miss parts of the target objects. For example, the front view only captures the cupboard (highlighted with a red circle), while the left and right shoulder views only show the sugar (already grasped by the end-effector and highlighted with green circles). These incomplete observations may lead to failed operations. In contrast, our proposed TAVP is designed to dynamically explore and re-render informative viewpoints that maximize coverage of target-relevant information, thereby improving the reliability of manipulation outcomes.
@@
Beyond acquiring complete and informative visual features, a scalable and task-sensitive visual encoder is also critical for accurate perception in multi-task settings. Although prior language-conditioned transformer-based methods, such as RVT [155] and RVT-2 [143], enable multi-task capabilities by attending to instruction tokens, their shared encoders still suffer from task interference, especially when handling visually and semantically diverse tasks (e.g., pick an apple vs. open a drawer). This entanglement limits the generalization and scalability of such models across a broad range of manipulation behaviors.
@@
To address these challenges, we propose the Task-Aware View Planning (TAVP) model, which integrates viewpoint exploration with task-aware visual feature extraction for robust and generalizable multi-task robotic manipulation. TAVP actively explores informative viewpoints using a delicately designed efficient reinforcement learning–based exploration policy, while employing a scalable task-aware Mixture-of-Experts (MoE) visual encoder to extract task-specific features. This design enables the model to obtain more complete and discriminative visual information, thereby improving robustness and generalization in multi-task action prediction.
@@
Specifically, our task-aware Mixture-of-Experts module, termed TaskMoE, improves multi-task accuracy and generalization through two key designs: (1) a dynamic expert routing mechanism guided by fused instruction and scene cues, rather than relying solely on task identifiers; and (2) a decoupled gating strategy that enables parameter sharing among semantically similar tasks, while isolating routing for semantically diverse ones. Based on the extracted task-specific features, we introduce the Multi-Viewpoint Exploration Policy (MVEP) to train a viewpoint exploration policy that selects optimal camera poses for each task. This Policy predicts the optimal camera re-rendering viewpoints based on the global point cloud. It then renders corresponding 2D images from these viewpoints for subsequent feature extraction and action prediction. These features possess enhanced global awareness with richer information, capable of improving robotic performance.
@@
To comprehensively evaluate the effectiveness of our TAVP in enhancing robotic spatial understanding and action generation performance, as well as its multi-task perception and decision-making capabilities, we conducted rigorous benchmark testing on 18 diverse RLBench [19] simulation tasks covering multiple tasks and scenarios. Our contributions are summarized as follows:
@@
  • Multi-Viewpoint Exploration Policy (MVEP) : The MVEP enables robots to actively explore optimal observation viewpoints (dynamic multi-view re-rendering), effectively addressing occlusion and insufficient viewpoint problems, and enhancing 3D perception capabilities.
@@
  • Superior Robotic Manipulation Performance: Extensive experiments on 18 RLBench simulation tasks demonstrate that our TAVP significantly outperforms existing baselines in multi-task manipulation w.r.t accuracy and robustness.
Related Works
Multi-task Learning in Manipulation
@@
The core challenges for general-purpose robots lie in task generalization capabilities and multi-task execution in complex environments. In the past, multi-task learning [99, 100, 143, 101] has achieved remarkable progress. It primarily consists of two approaches: Modular solutions decompose skills into reusable basic units (e.g., perception, planning, and control modules) [95]. Though interpretable, they face the combinatorial explosion problem—as task complexity increases, the cost of designing interfaces and coordinating between modules grows exponentially, leading to system rigidity. For instance, ManipGen [98] consumed massive training resources to train over 1,000 task-specific experts. Conversely, end-to-end solutions employ a single dense model to directly map perception to action [38, 40, 41, 30, 42, 43, 97]. While simplifying the pipeline, significant differences in visual feature distributions and motion patterns across tasks (e.g., grasping fragile objects vs. tightening screws) cause the model to struggle with convergence during training due to parameter conflicts. The Mixture of Experts (MoE) [12] addresses the scaling bottleneck of large models through a sparse activation mechanism. SDP [14] integrated MoE into diffusion policies to address multi-task learning and task transfer. Differently, we propose a task-aware MoE (TaskMoE) to resolve feature conflicts between different robotic manipulation tasks, activating distinct task-specific experts for different tasks, thereby enhancing multi-task execution performance and generalization capabilities.
Vision-Language-Action Model
Recently, numerous VLA models [42, 38, 40, 143, 79] have been proposed. Unlike traditional approaches involving pose estimation [110, 112] and path planning [113, 114], this significantly streamlines the model pipeline. With the advancement of large language models (LLMs) and vision-language models (VLMs), these large models have been applied in robotics [39, 119, 73, 94, 74, 111], leveraging their powerful priors to provide world knowledge for robots, thereby enhancing model generalization and performance. However, most current VLA models rely on a single or fixed viewpoint, making it difficult for robots to achieve comprehensive perception of 3D space, which limits action prediction capabilities. To address this, we propose a perception scheme featuring dynamic multi-view rerendering to enhance the spatial perception and comprehension capabilities, consequently improving the robot’s execution performance in complex scenarios.
@@
Figure 2: The overview of our Task-Aware View Planning (TAVP) framework. The input of this framework is four RGB-D images from fixed viewpoints. First, it converts them into point clouds and aggregates them into a global point cloud in the world coordinate system. Then, it diverges into two branches: One branch (orange) performs Coarse Grounding (a) to predict the approximate position of the end-effector. Subsequently, it moves the center of the global point cloud to this predicted position, performs scaling and cropping, retaining the important point cloud region. Another branch (green) receives the global point cloud, passes it through MVEP to predict the optimal camera parameters for the observation viewpoint. Then, using these parameters, it renders a 2D image from the point cloud processed by the red branch. This rendered image is fed into the Fine Grounding (b) to predict the final robot action, including the end-effector position, rotation, gripper status, and collision state.
Reinforcement Learning in Robotics
@@
Reinforcement learning (RL) has achieved significant accomplishments in robotics [102]. However, implementing RL starting from scratch typically necessitates complex reward mechanism and training paradigm design [103, 108, 107, 105, 106]. Recently, several works [104, 88] have demonstrated promising results by using RL to fine-tune pre-trained models. To optimize our MVEPN (Multi-Viewpoint Exploration Policy Network), we designed pseudo-environment interactions and a view-aware reward function, enabling policy learning based on a pre-trained model using offline expert data. Furthermore, to adapt between the MVEPN and the action network, we introduce supervised post-training to coordinate the robot’s ability to “See and Act”, thereby enhancing its task execution performance.
Methodology
Overview
@@
Our proposed Task-Aware View Planning (TAVP) framework aims to identify optimal viewpoints for accurate and robust robotic manipulation, guided by task-specific visual feature extraction. The full pipeline is illustrated in Fig. 2. TAVP takes as input a language instruction, the current visual observations from RGB-D cameras, and the current gripper state. To enable the model to explore the optimal observations from arbitrary viewpoints, we first reconstruct a 3D point cloud of the scene using the input RGB-D images. To narrow down the search space and enable task-aligned view selection, we leverage the coarse prediction stage from RVT-2 [143] to generate an area of interest.However, unlike RVT-2, which extracts visual features for all tasks using a shared Multi-View Transformer (MVT), we introduce a Task-Aware Mixture-of-Experts (TaskMoE) module before MVT to route instructions to specialized expert encoders. This design enables more precise and task-aligned visual feature extraction, which benefits both viewpoint selection and action prediction. Starting from the identified area of interest, we employ a Multi-Viewpoint Exploration Policy Network (MVEPN) to search for optimal camera poses that maximize the visibility of the target object and the end-effector. The selected viewpoints are then re-rendered into image observations and processed using another TaskMoE-based MVT before being passed to the action prediction model. For action prediction, we also upgrade the autoregressive action sequence model proposed in ARP [79] with our proposed TaskMoE to attain more task-specific feature extraction and action prediction.
@@
In the following sections, we detail the design of: 1) the TaskMoE, 2) the MVEPN, 3) the training strategy.
TaskMoE
@@
To address the inherent heterogeneity of complex manipulation tasks in multi-task learning, where different tasks often require substantially distinct visual representations and action policies, we introduce a task-aware Mixture-of-Experts module (TaskMoE), as illustrated in Fig. 3. Compared to prior MoE-based approaches [12, 14], our TaskMoE introduces two key innovations.
@@
First, instead of relying solely on task identifiers for expert selection, we incorporate richer instruction- and scene-related cues to guide expert routing more effectively, which is crucial for accurate multi-task robotic manipulation. Specifically, as shown in Fig. 3, we design a cross-modality module that employs a cross-attention mechanism to model the interaction between instruction and visual information. The resulting context-aware features are then fused with the task identifier via a Feature-wise Linear Modulation (FiLM) layer [181], enabling more adaptive and task-sensitive expert selection.
@@
Multi-Viewpoint Exploration Policy
@@
In this work, we represent each camera pose using a look-at model [182], which decouples the camera position and orientation via spherical coordinates. Therefore, the explored views are parameterized as 5-dimensional vectors $\mathbf{p}^{i}=(\theta^{i},\phi^{i},r^{i},\theta_{\text{up}}^{i},\phi_{\text{up}}^{i})\in\mathbb{R}^{5}$, where $(\theta,\phi,r)$ specify the camera center in spherical coordinates relative to the origin, $(\theta_{\text{up}},\phi_{\text{up}})$ define the orientation of the up vector, and $i\in[0,K-1]$ means the $i_{th}$ view. So, all camera poses are $\mathbf{p}=\{\mathbf{p}^{i}|i=0,...,K-1\}$.
To facilitate gradient-based policy optimization using reinforcement learning, we model the camera pose parameters as samples from a Gaussian distribution. Instead of directly predicting the viewpoint $\mathbf{p}^{i}$, the network outputs the mean and log-standard deviation, enabling differentiable sampling via the reparameterization trick.
@@
Training Strategy
The entire training process of our Task-Aware View Planning (TAVP) model consists of three stages:
Stage 1: In the first stage, we train a fixed-view variant of TAVP using three default rendering perspectives: front, left, and top. This stage follows a similar training pipeline to RVT-2 [143], where the overall loss comprises several components. Specifically, $\mathcal{L}_{hc}$ and $\mathcal{L}_{hf}$ denote the cross-entropy losses over heatmaps produced by the coarse and fine grounding modules, respectively. The ground-truth heatmaps are generated using a truncated Gaussian distribution centered at the 2D projection of the ground-truth 3D location, as described in RVT [155]. Additionally, the loss includes the end-effector rotation loss $\mathcal{L}_{rot}$, computed as the cross-entropy loss for each of the Euler angles, as well as the gripper state loss $\mathcal{L}_{gri}$ and collision indicator loss $\mathcal{L}_{col}$, both formulated as binary classification losses. The total objective for the first stage is given by:

### v4

# Learning to See and Act: Task-Aware Virtual View Exploration for Robotic Manipulation

@@ Abstract
Recent vision-language-action (VLA) models for multi-task robotic manipulation commonly rely on static viewpoints and shared visual encoders, which limit 3D perception and cause task interference, hindering robustness and generalization. In this work, we propose Task-aware Virtual View Exploration (TVVE), a framework designed to overcome these challenges by integrating virtual view exploration with task-specific representation learning. TVVE employs an efficient exploration policy, accelerated by a novel pseudo-environment, to acquire informative views. Furthermore, we introduce a Task-aware Mixture-of-Experts (TaskMoE) visual encoder to disentangle features across different tasks, boosting both representation fidelity and task generalization. By learning to see the world in a task-aware way, TVVE generates more complete and discriminative visual representations, demonstrating significantly enhanced action prediction across a wide array of manipulation challenges. To further validate the robustness and generalization capability of TVVE under out-of-distribution (OOD) settings, we construct a challenging benchmark, RLBench-OG, covering various visual perturbations and camera pose variations. Extensive experiments on RLBench and RLBench-OG show that our TVVE achieves superior performance over state-of-the-art approaches. In real-robot experiments, TVVE demonstrates exceptional performance and generalizes robustly in multiple OOD settings, including visual disturbances and unseen instructions. Visual results and code are provided at: TAVP.
1 Introduction
A general-purpose robotic system is expected to accurately perform numerous manipulation tasks [liu2025aligning]. With the development of vision-language-action (VLA) models [kimopenvla, rt-1, rt-2, livision, cheang2024gr, bu2024towards, li2025bridgevla, bu2025univla, cen2025worldvla, liu2025hybridvla, team2024octo, lin2025onetwovla] and transformer-based architectures [ACT, rt-1, rt-2, hou2024diffusion, chi2023diffusion], many end-to-end robotic manipulation frameworks (e.g., OpenVLA[kimopenvla], $\pi_{0}$[black2410pi0]) have achieved promising results on complex and fine-grained tasks. Typically, these models rely on observations captured from a single view [kimopenvla, chi2023diffusion, shridhar2022cliport] or a few fixed viewpoints [ke20243ddiffactor, shridhar2023perceiver] to guide action prediction. While this may suffice in simple environments with only a few objects, it becomes problematic in cluttered or dynamic scenes—particularly when objects or the end-effector are repositioned during task execution.
Figure 1: Motivation Illustration. Observations captured from fixed cameras often miss parts of the target objects. For example, the front view only captures the cupboard (highlighted with a red circles), while the left and right shoulder views only show the sugar (already grasped by the end-effector and highlighted with green circles). These incomplete observations may lead to failed operations. In contrast, our proposed TVVE is designed to dynamically explore and re-render informative viewpoints that maximize coverage of target-relevant information, thereby improving the reliability of manipulation outcomes.
@@ 1 Introduction
Beyond acquiring complete and informative visual features, a scalable and task-sensitive visual encoder is also critical for accurate perception in multi-task settings. Although prior language-conditioned transformer-based methods, such as RVT [goyal2023rvt] and RVT-2 [goyal2024rvt2], enable multi-task capabilities by attending to instruction tokens, their shared encoders still suffer from task interference, especially when handling visually and semantically diverse tasks (e.g., pick an apple vs. open a drawer). This entanglement limits the generalization and scalability of such models across a broad range of manipulation behaviors.
@@
To address these challenges, we propose the Task-aware Virtual View Exploration (TVVE) model, which integrates viewpoint exploration with task-aware visual feature extraction for robust and generalizable multi-task robotic manipulation. TVVE explores informative viewpoints using a delicately designed efficient reinforcement learning–based exploration policy, while employing a scalable task-aware Mixture-of-Experts (MoE) visual encoder to extract task-specific features. This design enables the model to obtain more complete and discriminative visual information, thereby improving robustness and generalization in multi-task action prediction.
@@
Specifically, our task-aware Mixture-of-Experts module, termed TaskMoE, improves multi-task accuracy and generalization through two key designs: (1) a dynamic expert routing mechanism guided by fused instruction and scene cues, rather than relying solely on task identifiers; and (2) a decoupled gating strategy that enables parameter sharing among semantically similar tasks, while isolating routing for semantically diverse ones. Based on the extracted task-specific features, we introduce the Multi-Viewpoint Exploration Policy (MVEP) that adaptively selects task-specific camera poses based on the global point cloud. By re-rendering 2D observations from the explored viewpoints, MVEP enriches visual perception with globally aware features, leading to more informed action prediction and improved manipulation success.
@@
To comprehensively evaluate the effectiveness of our TVVE in enhancing robotic spatial understanding and action generation performance, as well as its multi-task perception and decision-making capabilities, we conducted rigorous benchmark testing on 18 diverse RLBench [james2020rlbench] simulation tasks covering multiple tasks and scenarios. Our contributions are summarized as follows:
@@
  • Multi-Viewpoint Exploration Policy (MVEP) : The MVEP enables robots to explore optimal virtual observation viewpoints (dynamic multi-view re-rendering), effectively addressing occlusion and insufficient viewpoint problems, and enhancing 3D perception capabilities.
@@
  • RLBench-OG: We construct a new benchmark to comprehensively evaluate the model robustness and generalization under out-of-distribution (OOD) settings, covering visual perturbations such as occlusions, background changes, lighting variations, texture shifts, and camera pose changes.
@@
  • Superior Robotic Manipulation Performance: Extensive experiments on RLBench, RLBench-OG and real-robot tasks demonstrate that our TVVE significantly outperforms existing baselines in multi-task manipulation w.r.t accuracy and robustness.
2 Related Works
2.1 Multi-task Learning in Manipulation
@@
The core challenges for general-purpose robots lie in task generalization capabilities and multi-task execution in complex environments. In the past, multi-task learning [deisenroth2014multi, ze2023gnfactor, goyal2024rvt2, ma2024hierarchical] has achieved remarkable progress. It primarily consists of two approaches: Modular solutions decompose skills into reusable basic units (e.g., perception, planning, and control modules) [mao2024robomatrix]. Though interpretable, they face the combinatorial explosion problem—as task complexity increases, the cost of designing interfaces and coordinating between modules grows exponentially, leading to system rigidity. For instance, ManipGen [dalallocal] consumed massive training resources to train over 1,000 task-specific experts. Conversely, end-to-end solutions employ a single dense model to directly map perception to action [team2024octo, ze20243d, wang2024poco, chi2023diffusion, shridhar2022cliport, nair2022learning, haldar2024baku]. While simplifying the pipeline, significant differences in visual feature distributions and motion patterns across tasks (e.g., grasping fragile objects vs. tightening screws) cause the model to struggle with convergence during training due to parameter conflicts. The Mixture of Experts (MoE) [shazeer2017outrageously] addresses the scaling bottleneck of large models through a sparse activation mechanism. SDP [wangsparse] integrated MoE into diffusion policies to address multi-task learning and task transfer. Differently, we propose a task-aware MoE (TaskMoE) to resolve feature conflicts between different robotic manipulation tasks, activating distinct task-specific experts for different tasks, thereby enhancing multi-task execution performance and generalization capabilities.
@@
Figure 2: The overview of our Task-aware Virtual View Exploration (TVVE) framework. The input of this framework is multiple RGB-D images from fixed viewpoints. First, it converts them into point clouds and aggregates them into a global point cloud in the world coordinate system. Then, it diverges into two branches: One branch (orange) performs Coarse Grounding to predict the approximate position of the end-effector. Subsequently, it moves the center of the global point cloud to this predicted position, performs scaling and cropping, retaining the important point cloud region. Another branch (green) receives the global point cloud, passes it through MVEP to predict the optimal camera parameters for the observation viewpoint. Then, using these parameters, it renders a 2D image from the point cloud processed by the red branch. This rendered image is fed into the Fine Grounding to predict the final robot action, including the end-effector position, rotation, gripper status, and collision state.
2.2 Vision-Language-Action Model
Recently, numerous VLA models [shridhar2022cliport, team2024octo, ze20243d, goyal2024rvt2, zhang2025autoregressive] have been proposed. In contrast to traditional pipelines that rely on explicit pose estimation [fang2020graspnet, deng2020self] and path planning [prianto2020path, zhao2025deep], VLA models directly map linguistic instructions and visual observations to low-level actions, substantially simplifying the control stack. With the advancement of large language models (LLMs) and vision-language models (VLMs), these large models have been applied in robotics [kimopenvla, rt-1, rt-2, livision, cheang2024gr, bu2024towards], leveraging their powerful priors to provide world knowledge for robots, thereby enhancing model generalization and performance. However, most existing VLA models operate on a single or fixed viewpoint, limiting a robot’s ability to form comprehensive 3D spatial understanding and thereby constraining action prediction. To overcome this limitation, we introduce a dynamic multi-view re-rendering perception scheme that enhances spatial perception and scene comprehension, ultimately improving execution robustness in complex environments.
2.3 Reinforcement Learning in Robotics
@@
Reinforcement learning (RL) has achieved significant accomplishments in robotics [tang2025deep]. However, implementing RL starting from scratch typically necessitates complex reward mechanism and training paradigm design [hu2024flare, khetarpal2022toward, luo2024serl, rajeswaran2017learning, uchendu2023jump]. Recently, several works [guo2025improving, lu2025vla] have demonstrated promising results by using RL to fine-tune pre-trained models. To optimize our MVEP (Multi-Viewpoint Exploration Policy), we designed pseudo-environment interactions and a view-aware reward function, enabling policy learning based on a pre-trained model using offline expert data. Furthermore, to adapt between the MVEP and the action policy, we introduce supervised post-training to coordinate the robot’s ability to “See and Act”, thereby enhancing its task execution performance.
3 Methodology
3.1 Overview
@@
Our proposed Task-aware Virtual View Exploration (TVVE) framework aims to identify optimal viewpoints for accurate and robust robotic manipulation, guided by task-specific visual feature extraction. The full pipeline is illustrated in Fig. 2. TVVE takes as input a language instruction, the current visual observations from RGB-D cameras, and the current gripper state. To enable the model to explore the optimal observations from arbitrary viewpoints, we first reconstruct a 3D point cloud of the scene using the input RGB-D images. To narrow down the search space and enable task-aligned view selection, we leverage the coarse prediction stage from RVT-2 [goyal2024rvt2] to generate an area of interest. However, unlike RVT-2, which extracts visual features for all tasks using a shared Multi-View Transformer (MVT), we integrate a Task-Aware Mixture-of-Experts (TaskMoE) module within the MVT to route instructions to specialized expert encoders. This design enables more precise and task-aligned visual feature extraction, which benefits both viewpoint selection and action prediction. Starting from the identified area of interest, we employ a Multi-Viewpoint Exploration Policy (MVEP) to search for optimal camera poses that maximize the visibility of the target object and the end-effector. The selected viewpoints are then re-rendered into image observations and processed using another TaskMoE-based MVT before being passed to the action prediction model. For action prediction, we also upgrade the autoregressive action policy proposed in ARP [zhang2025autoregressive] with our proposed TaskMoE to attain more task-specific feature extraction and action prediction.
@@
In the following sections, we detail the design of: 1) the TaskMoE, 2) the MVEP, and 3) the training strategy.
3.2 TaskMoE
@@
To address the inherent heterogeneity of complex manipulation tasks in multi-task learning, where different tasks often require substantially distinct visual representations and action policies, we introduce a task-aware Mixture-of-Experts module (TaskMoE), as illustrated in Fig. 3. Compared to prior MoE-based approaches [shazeer2017outrageously, wangsparse], our TaskMoE introduces two key innovations.
@@
First, instead of relying solely on task identifiers for expert selection, we incorporate richer instruction- and scene-related cues to guide expert routing more effectively, which is crucial for accurate multi-task robotic manipulation. Specifically, as shown in Fig. 3, we design a cross-modality module that employs a cross-attention mechanism to model the interaction between instruction and visual information. The resulting context-aware features are then fused with the task identifier via a Feature-wise Linear Modulation (FiLM) layer [Film], enabling more adaptive and task-sensitive expert selection.
@@
3.3 Multi-Viewpoint Exploration Policy
@@
In this work, we represent each camera pose using a look-at model [lookatmodel], which decouples the camera position and orientation via spherical coordinates. Therefore, the explored views are parameterized as 5-dimensional vectors $\mathbf{p}^{i}=(\theta^{i},\phi^{i},r^{i},\theta_{\text{up}}^{i},\phi_{\text{up}}^{i})\in\mathbb{R}^{5}$, where $(\theta,\phi,r)$ specify the camera center in spherical coordinates relative to the origin, $(\theta_{\text{up}},\phi_{\text{up}})$ define the orientation of the up vector, and $i\in[0,K-1]$ means the $i_{th}$ view. So, all camera poses are $\mathbf{p}=\{\mathbf{p}^{i}|i=0,...,K-1\}$. Additionally, the camera is perpetually oriented toward the coordinate origin $O(0,0,0)$.
@@
3.4 Training Strategy
The entire training process of our Task-aware Virtual View Exploration (TVVE) model consists of three stages:
Stage 1: In the first stage, we train a fixed-view variant of TVVE using three default rendering perspectives: front, left, and top. This stage follows a similar training pipeline to RVT-2 [goyal2024rvt2], where the overall loss comprises several components. Specifically, $\mathcal{L}_{hc}$ and $\mathcal{L}_{hf}$ denote the cross-entropy losses over heatmaps produced by the coarse and fine grounding modules, respectively. The ground-truth heatmaps are generated using a truncated Gaussian distribution centered at the 2D projection of the ground-truth 3D location, as described in RVT [goyal2023rvt]. Additionally, the loss includes the end-effector rotation loss $\mathcal{L}_{rot}$, computed as the cross-entropy loss for each of the Euler angles, as well as the gripper state loss $\mathcal{L}_{gri}$ and collision indicator loss $\mathcal{L}_{col}$, both formulated as binary classification losses. The total objective for the first stage is given by: